OpenStudy (anonymous):

Which of the following equations is the equation 2 - 3y + 6x = 0 written in slope-intercept form? y = -2x - 2/3 y = 1/2x + 2 y = 2x + 2/3

OpenStudy (anonymous):

well slope intercept form is y = mx + b

OpenStudy (anonymous):

the key is to get y by itself in order to do that, we must move things over 2 - 3y + 6x = 0 the first thing im going to do is subtract 6x on both sides 2-3y+6x=0 -6x -6x ------------------------ 2-3y= -6x Next we will subtract 2 on both sides -2 -2 ----------------------- -3y=-6x-2 Now were going to divide bothe sides by -3 the -3 cancels on this side --> -3y/-3= -6x/-3-2/-3 y= 2x+2/3 So the answer is the last one C

OpenStudy (anonymous):

@aussy123 Please don't give answers :/ openstudy.com/code-of-conduct

OpenStudy (anonymous):

Im sorry I was trying to explain in a way he/she would understand